# Eaman Jahani

### v4

@@ A.2.2 Target Images
We randomly assigned participants to 1 of 15 target images. The set of target images consisted of 5 images each from 3 different broad categories: business and marketing, graphic design, and architectural photography. We chose these to represent the use cases suggested by the prompt categories on https://promptbase.com/, a leading marketplace for image generation prompts. The images vary in color, style, content, and complexity within and across categories. These images can be found online linked to the pre-registration document: https://osf.io/ejbtp. As we discuss in Section F, performance, and variability of performance varied substantially across images. In other words, some images were much easier than others to replicate with the generative models, which we view as additional evidence that the set of 15 images was reasonably diverse.
@@ A.3 Subjects
Our Prolific-recruited US sample (N = 2,059) was limited to fluent English speakers, and we prevented participants from completing the task more than once. We also prevented users from completing the task on mobile devices or tablets. Data was collected between December 12, 2023 and December 19, 2023. Participants were guaranteed a payment of \$4 USD for completing the task and could earn an additional \$8 USD (a 200% bonus) if they ranked in the top 20% of participants in DreamSim of their image most similar to the target (construction of DreamSim is described in section B.5.1). The median time to complete our entire task, including a demographic survey, was 22 minutes. Given that 20% of subjects received a bonus, the average compensation for participants in our study was \$5.60 USD per person, or about \$15 USD per hour. We explained the payment and incentive scheme to participants in full multiple times during the onboarding phase of the experiment, and asked participants to confirm their understanding before they were allowed to complete the task. The onboarding process also included multiple attention checks; participants who failed the first check were immediately disqualified. For subsequent checks, participants were required to retry until they demonstrated understanding.
@@ A.4 Model Endpoints
Appendix B Measurement and Variables
B.1 Survey Data
For each participant, we used a Qualtrics survey to collect demographic information, information on the participant’s skills that may be relevant to generative AI use, and information on the participant’s attitude towards generative AI. This data includes:
@@ B.1 Survey Data
B.2 Prompt Data
For each prompt, we record the text of the participant’s prompt, the order in which it was submitted, the timestamp of submission, and for the DALL-E 3 treatment arms, the revised prompt returned by the model.
B.3 Image Data
For each prompt, the following images were collected:
@@ B.3 Image Data
    Post-hoc resampled images (Azure OpenAI endpoint): For any given prompt, the output of the text-to-image model is stochastic. To better approximate the expected image from a given prompt, we generated 20 additional images for each prompt after the experiment concluded. We provide full details on this procedure in Section C. These images were generated from December 26, 2023 - January 27, 2024. These images are not used for analyses presented in the main text, but were used for other pre-registered analyses. These additional analyses are discussed in Section F.
@@
B.4 Sample Construction
The sample we analyze in the main text of our paper is constructed using the process described as follows.
@@ B.4 Sample Construction
    B.4.1 “Off-Topic” Robustness Check
@@
B.5 Dependent Variables
B.5.1 Image Similarity
We pre-registered two quantitative measures of image similarity: the cosine similarity of CLIP embedding vectors and a recently developed measure called ‘DreamSim’ (Fu et al., 2023). In the main text, we present analyses using CLIP embedding cosine similarity, since it is likely more familiar to readers. Our results are qualitatively and quantitatively similar using DreamSim instead.
  • CLIP Embedding Cosine Similarity: To calculate CLIP embedding cosine similarity, we first generated CLIP embedding vectors (Radford et al., 2021) from Hugging Face (CLIP-embeddings) for each participant-generated image and for each target image. Unlike traditional image embeddings that only encode visual features, CLIP embeddings also capture semantic relationships between images and descriptive text. We then calculated the cosine similarity between each participant-generated image’s CLIP embedding and the relevant target image’s CLIP embedding.
  • DreamSim: DreamSim is an image similarity measure proposed recently by (Fu et al., 2023). The authors claim that relative to a measure such as CLIP embedding cosine similarity, DreamSim measures image similarity in a way that more effectively captures human visual perceptions of similarity. Because the original DreamSim metric outputs a distance measure, we invert this score $\tilde{D}=1-(\text{original DreamSim})$ to recast it as a similarity score. After doing so, both the inverted DreamSim and CLIP embedding cosine similarity are closer to 1 when two images are more similar and closer to 0 when two images are more dissimilar.
We find that these two measures of image similarity are highly correlated in our sample ($\rho_{pearson}=0.763$, 95% CI: [0.755 0.770]), and our main results are robust to the use of either measure. We present the results obtained when conducting our main text analyses using DreamSim in Section E.E.1.
B.5.2 Prompt Length
@@
B.5.3 Embedding-based Prompt Similarity
We calculate two measures of embedding-based prompt similarity: successive similarity and aggregate similarity. Both measures use the vector embedding representation of each prompt in our sample, which we obtained using OpenAI’s text-embedding-3-small model (Neelakantan et al., 2022). The two similarity measures are defined as follows:
@@ B.5.3 Embedding-based Prompt Similarity
B.5.4 Successive Prompt Token Sort Ratio
Starting with each participant’s second prompt, we also calculated the token sort ratio (TSR) of each prompt $p_{i,n}$ to the immediately preceding prompt $p_{i,n-1}$. TSR is a fuzzy string-matching technique (Singla and Garg, 2012) that provides a continuous measure of how similar two strings are. We refer the reader to (Bosker, 2021) for a more in-depth description of how TSR is calculated.
B.5.5 Successive Prompt ‘Contains Previous Prompt’ Dummy
@@
B.5.6 Prompt Composition
We use the spaCy v3.7.4 Python package’s en_core_web_sm model to tag the parts of speech (POS) in each prompt. SpaCy’s models utilize the ”universal POS tags” from the Universal Dependencies framework for grammar annotation Universal POS tags (2024). These tags encompass parts of speech such as adjectives, adverbs, nouns, and verbs. The model tags each word in a prompt according to this framework, after which we count the total number of words corresponding to each part of speech for each prompt.
B.5.7 Strategic Shifts
In addition to calculating the successive and aggregate similarity of prompts written by particular users, we also attempt to identify particular moments when participants shift their approach to prompting. In order to do so, we adapt a method proposed in (Torricelli et al., 2023) (because they are conducting research in a different context, (Torricelli et al., 2023) refer to these shifts as “topical transitions” as opposed to “strategic shifts”). To identify these strategic shifts, we first calculate the mean cosine similarity ($MCS$) for the embedding vectors of every possible pair of prompts submitted in response to a given target image, $t$:
@@ B.5.7 Strategic Shifts
$SS(\mathbf{p}_{i,n,t})=\begin{cases}1&\text{if }\textrm{CosineSim}(\mathbf{E(p_{i,n,t})},\mathbf{E(p_{i-1,n,t})})<MCS_{t}\\ 0&\text{otherwise}\end{cases}$ (4)
@@
Appendix C Methods
C.1 Stratification
The results shown in the main text and in the SI are mostly stratified by reference image and iteration. In some analyses, we have stratified only on the reference image (e.g., for analyses presented at the iteration level). The exact stratification for each finding is indicated in section D. To stratify our results, we take a weighted average across $j=1,...,J$ cells defined by our stratification variables.
@@ C.1 Stratification
To calculate the variance (and standard error) of this sample mean we apply the following:
@@ C.2 Z-Scoring
We find statistically significant evidence for differences in the variability of performance across the 15 target images used in our experiments, which we discuss in Section F. In the main text, we also showed that performance increases across the attempts. To test whether our results are in some way due to this image-level or attempt-level variation, we replicate all analyses using the within-image-attempt Z-score of CLIP-cosine similarity of each image produced by participants in our experiment. Formally, this is:
$Z(\textrm{CosineSim}_{i,n,t})=\frac{\textrm{CosineSim}_{i,n,t}-\textrm{Mean}_{n,t}\left(\textrm{CosineSim}_{i,n,t}\right)}{\textrm{SD}_{n,t}\left(\textrm{CosineSim}_{i,n,t}\right)},$ (5)
@@
where $\textrm{CosineSim}_{i,n,t}$ is the cosine similarity of user $i$’s image in attempt $n$ to target image $t$. The mean and standard deviation are computed per each image-attempt, but over the DALL-E 2 and DALL-E 3 treatment arms. We also use the rescaling above to test the robustness of our DreamSim-based analyses. Almost all the robustness analysis reported here use the Z-score scaled measure of performance within each image-attempt set. The only exception is any analysis that examines improvements and prompts across attempts, e.g. figure 1B, where Z-scores are computed only within each target image, and across all attempts of that image.
C.3 Accounting for Model Stochasticity
@@
The output that generative AI models return in response to a given prompt is stochastic. The strength of this stochasticity is controlled by a model parameter referred to as the temperature, which could not be edited using the DALL-E API at the time of our experiment. To account for this model stochasticity, we generated 10 images for each prompt submitted by participants for all arms. We were then able to calculate the similarity between each replication and its corresponding target image, and calculate an “expected” CLIP cosine similarity and DreamSim score for each prompt by averaging over these samples. Given the replicated images, we can also calculate the standard deviation of cosine similarity induced by this stochasticity. With the expected cosine similarity and its standard deviation per prompt, we can compute a normalized Z-score for the observed image relative to its replication distribution. This Z-score measures the extent to which the observed image is better or worse than what’s expected for that prompt and will be used for further analysis in section F.
@@
We generated these additional samples for both the original prompts on their assigned treatment arms, as well as replaying on the counterfactual arms, as introduced in Figure 2 in the main text. Importantly, OpenAI updated its content filters between our initial experiment and image re-sampling. As a result, some prompts that originally produced images either generated no images or fewer images than requested during our regeneration attempts. This affected 1.8% (371 out of 18,990 prompts) of the data in our sample under the ”replaying” procedure (Section B.B.3.3).
Appendix D Main Text Analyses
D.1 Task Performance and ATEs

### v5

@@ A.2.2 Target Images
We randomly assigned participants to 1 of 15 target images. The set of target images consisted of 5 images each from 3 different broad categories: business and marketing, graphic design, and architectural photography. We chose these to represent the use cases suggested by the prompt categories on https://promptbase.com/, a leading marketplace for image generation prompts. The images vary in color, style, content, and complexity within and across categories. These images can be found online linked to the pre-registration document: URL removed to maintain anonymity. Performance, and variability of performance varied substantially across images. In other words, some images were much easier than others to replicate with the generative models, which we view as additional evidence that the set of 15 images was reasonably diverse.
@@ A.3 Subjects
Our Prolific-recruited US sample (N = 2,059) was limited to fluent English speakers, and we prevented participants from completing the task more than once. We also prevented users from completing the task on mobile devices or tablets. Data was collected between December 12, 2023 and December 19, 2023. Participants were guaranteed a payment of \$4 USD for completing the task and could earn an additional \$8 USD (a 200% bonus) if they ranked in the top 20% of participants in DreamSim of their image most similar to the target (construction of DreamSim is described in section C.5.1). The median time to complete our entire task, including a demographic survey, was 22 minutes. Given that 20% of subjects received a bonus, the average compensation for participants in our study was \$5.60 USD per person, or about \$15 USD per hour. We explained the payment and incentive scheme to participants in full multiple times during the onboarding phase of the experiment, and asked participants to confirm their understanding before they were allowed to complete the task. The onboarding process also included multiple attention checks; participants who failed the first check were immediately disqualified. For subsequent checks, participants were required to retry until they demonstrated understanding.
@@ A.4 Model Endpoints
A.5 Image Similarity Metrics
The primary outcome in our experiment is the similarity between each participant-generated image and the assigned target image, measured using the cosine similarity of CLIP embeddings (Radford et al. 2021). CLIP (Contrastive Language–Image Pretraining) is a neural network trained to jointly embed images and text into a shared latent space, such that semantically or visually similar items lie close together. By embedding both the target image and each generated image into this space and computing the cosine similarity between them, we obtain a quantitative measure of how closely the generated image matches the target along both visual and conceptual dimensions.
Because the output of each generative model is stochastic, the same prompt can yield different images across attempts. To account for this variability, we generated 10 images for each prompt and computed their cosine similarity to the target image individually. We then averaged these 10 similarity scores to produce an expected quality score for each prompt—the primary outcome variable used in our analysis. As a robustness check, we replicated all analyses using DreamSim, a recently developed perceptual similarity metric that aligns more closely with human judgments. [^9] DreamSim (Fu et al. 2023) is designed to better capture human perceptions of image similarity than traditional embedding-based methods. Our results are robust to using DreamSim in place of CLIP cosine similarity. Full DreamSim-based analyses can be found in Appendix F. The two measures were highly correlated, and our findings were consistent across both.
A.6 Pre-registration
This study was pre-registered. The pre-registration document included our hypotheses, planned analyses, and sample size justification. The pre-registration document can be found at URL removed to maintain anonymity.
Appendix B Example Prompts
Table 2 presents the complete set of prompts and their corresponding generated images that were used in our analysis shown in Figure 1. The table displays three example target images (woman with shopping basket, abstract painting with cross, and online shopping interface), with three different participant-generated versions of each. For each target, we show examples of varying quality, arranged from most similar to least similar. The prompts are presented verbatim from participant submissions, preserving all original formatting and punctuation (or lack thereof).
Table 2: Participant prompts for the images provided in Figure 1
Image Prompt
[Uncaptioned image] “create an image of a woman with blond hair, wearing a surgical mask, black gloves, a white t-shirt, and holding a red shopping basket full of groceries”
[Uncaptioned image] “A lady has long blonde hair, wears BLACK COLORED GLOVES and a BLUE COLORED mask, with a white blouse on. She is holding a RED PLASTIC BASKET filled with different kinds of fruit.”
[Uncaptioned image] “Generate a stock photo of a woman holding a basket of groceries during the pandemic. Make sure she has on black gloves.”
[Uncaptioned image] “Thank you! Please create a mostly white painting with red and blue lines in the center, shaped a bit like the # symbol. Add red and blue splatter as well, but not too much”
[Uncaptioned image] “Water color style. Japanese vibes. White background. Square outline in the middle of the page is standing on its side at a 12 degree angle. The inside of the square is white. Two legs are blue and two are red. T”
[Uncaptioned image] “painting of two x’s intersecting to make a slanted square. The square is at a 35 degree angle with the left upper side being two red lines making the x and the lower right side two blue lines. The painting only has red and blue and the square with 2 x’s sits in the middle of the painting with light speckles of red and blue drops of paint scattered throughout the painting”
[Uncaptioned image] “pastels, abstract, bearded man on left in a pink shirt with blue sleeves pointing to a shirt box image, woman with pink hair in a blue dress and high heels holding the handle of a shopping cart full of purses on right side, pink and white awning in center top, 4 out of 5 stars on left side in pink, animation style. single pink dollar sign on the left. white search bar in the center with a magnifying glass icon on the right side of the search bar.”
[Uncaptioned image] “cartoon-y virtual self checkout. man shopping for clothes. woman pushing cart. phone that looks like a storefront in background.”
[Uncaptioned image] “at the store a man using a touch screen monitor to shop for a shirt and a woman walking by with a store cart full of gift”
Appendix C Measurement and Variables
This appendix provides detailed information about the data collection, measurement, and variables used in our analyses, expanding on the overview provided in the main text.
C.1 Survey Data
We collected additional participant information via a Qualtrics survey that included:
@@ C.1 Survey Data
C.2 Prompt Data
For each prompt, we recorded the text of the participant’s prompt, the order in which it was submitted, the timestamp of submission, and for the DALL-E 3 treatment arms, the revised prompt returned by the model.
C.3 Image Data
We collected three different sets of images:
@@ C.3 Image Data
    Post-hoc resampled images (Azure OpenAI endpoint): For any given prompt, the output of the text-to-image model is stochastic. To better approximate the expected image from a given prompt, we generated 20 additional images for each prompt after the experiment concluded. We provide full details on this procedure in Section D. These images were generated from December 26, 2023 - January 27, 2024. These images are not used for analyses presented in the main text, but were used for other pre-registered analyses. These additional analyses are discussed in Section G.
@@
C.4 Sample Construction
The sample analyzed in the main text was constructed as follows:
@@ C.4 Sample Construction
C.4.1 “Off-Topic” Robustness Check
@@
C.5 Dependent Variables
C.5.1 Image Similarity
We pre-registered two quantitative measures of image similarity: the cosine similarity of CLIP embedding vectors and a recently developed measure called ‘DreamSim’ (Fu et al. 2023). In the main text, we present analyses using CLIP embedding cosine similarity, since it is likely more familiar to readers. Our results are qualitatively and quantitatively similar using DreamSim instead.
  • CLIP Embedding Cosine Similarity: To calculate CLIP embedding cosine similarity, we first generated CLIP embedding vectors (Radford et al. 2021) from Hugging Face (OpenAI 2024) for each participant-generated image and for each target image. Unlike traditional image embeddings that only encode visual features, CLIP embeddings also capture semantic relationships between images and descriptive text. We then calculated the cosine similarity between each participant-generated image’s CLIP embedding and the relevant target image’s CLIP embedding.
  • DreamSim: DreamSim is an image similarity measure proposed recently by (Fu et al. 2023). The authors claim that relative to a measure such as CLIP embedding cosine similarity, DreamSim measures image similarity in a way that more effectively captures human visual perceptions of similarity. Because the original DreamSim metric outputs a distance measure, we invert this score $\tilde{D}=1-(\text{original DreamSim})$ to recast it as a similarity score. After doing so, both the inverted DreamSim and CLIP embedding cosine similarity are closer to 1 when two images are more similar and closer to 0 when two images are more dissimilar.
We find that these two measures of image similarity are highly correlated in our sample ($\rho_{pearson}=0.763$, 95% CI: [0.755 0.770]), and our main results are robust to the use of either measure. We present the results obtained when conducting our main text analyses using DreamSim in Section F.F.1.
C.5.2 Prompt Length
@@
C.5.3 Embedding-based Prompt Similarity
We calculate two measures of embedding-based prompt similarity: successive similarity and aggregate similarity. Both measures use the vector embedding representation of each prompt in our sample, which we obtained using OpenAI’s text-embedding-3-small model (Neelakantan et al. 2022). The two similarity measures are defined as follows:
@@ C.5.3 Embedding-based Prompt Similarity
C.5.4 Successive Prompt Token Sort Ratio
Starting with each participant’s second prompt, we also calculated the token sort ratio (TSR) of each prompt $p_{i,n}$ to the immediately preceding prompt $p_{i,n-1}$. TSR is a fuzzy string-matching technique (Singla and Garg 2012) that provides a continuous measure of how similar two strings are.
C.5.5 Successive Prompt ‘Contains Previous Prompt’ Dummy
@@
C.5.6 Prompt Composition
We use the spaCy v3.7.4 Python package’s en_core_web_sm model to tag the parts of speech (POS) in each prompt. SpaCy’s models utilize the ”universal POS tags” from the Universal Dependencies framework for grammar annotation Universal Dependencies Project (2024). These tags encompass parts of speech such as adjectives, adverbs, nouns, and verbs. The model tags each word in a prompt according to this framework, after which we count the total number of words corresponding to each part of speech for each prompt.
C.5.7 Strategic Shifts
In addition to calculating the successive and aggregate similarity of prompts written by particular users, we also attempt to identify particular moments when participants shift their approach to prompting. In order to do so, we adapt a method proposed in (Torricelli et al. 2023) (because they are conducting research in a different context, (Torricelli et al. 2023) refer to these shifts as “topical transitions” as opposed to “strategic shifts”). To identify these strategic shifts, we first calculate the mean cosine similarity ($MCS$) for the embedding vectors of every possible pair of prompts submitted in response to a given target image, $t$:
@@ C.5.7 Strategic Shifts
$SS(\mathbf{p}_{i,n,t})=\begin{cases}1&\text{if }\textrm{CosineSim}\big{(}\mathbf{E(p}_{i,n,t}),\mathbf{E(p}_{i-1,n,t})\big{)}\\ &\quad<MCS_{t}\\ 0&\text{otherwise}\end{cases}$ (9)
@@
Appendix D Methods
This appendix provides additional methodological details to supplement the analyses presented in the main text.
D.1 Stratification
The results shown in the main text and in the supplementary information are typically stratified by reference image and iteration. In some analyses, we have stratified only on the reference image (e.g., for analyses presented at the iteration level). The exact stratification for each finding is indicated in section E. To stratify our results, we calculate a weighted average across $j=1,...,J$ cells defined by our stratification variables:
@@ D.1 Stratification
To calculate the variance (and standard error) of this sample mean, we apply the following:
@@ D.2 Z-Scoring
Our analysis found statistically significant differences in performance variability across the 15 target images used in our experiments, as discussed in Section G. The main text also demonstrated that performance increases across successive attempts. To ensure our results are not driven by this image-level or attempt-level variation, we replicated all analyses using the within-image-attempt Z-score of CLIP-cosine similarity for each image produced by participants. Formally, this is:
$Z(\text{Sim}_{i,n,t})=\frac{\text{Sim}_{i,n,t}-\text{Mean}_{n,t}\left(\text{Sim}_{i,n,t}\right)}{\text{SD}_{n,t}\left(\text{Sim}_{i,n,t}\right)}$ (10)
@@
where $\textrm{Sim}_{i,n,t}$ is the cosine similarity of user $i$’s image in attempt $n$ to target image $t$. The mean and standard deviation are computed for each image-attempt pair, but across both the DALL-E 2 and DALL-E 3 treatment arms. We also applied this rescaling to test the robustness of our DreamSim-based analyses.
Nearly all robustness analyses reported here use the Z-score scaled measure of performance within each image-attempt set. The only exception is analyses that examine improvements and prompts across attempts (e.g., Figure 1B), where Z-scores are computed only within each target image and across all attempts for that image.
D.3 Accounting for Model Stochasticity
@@
Generative AI models produce stochastic outputs in response to a given prompt. This stochasticity is controlled by a parameter called temperature, which could not be modified using the DALL-E API at the time of our experiment. To account for this model stochasticity, we generated 10 images for each prompt submitted by participants across all treatment arms. We then calculated the similarity between each replication and its corresponding target image, and computed an ”expected” CLIP cosine similarity and DreamSim score for each prompt by averaging across these samples.
Using these replicated images, we also calculated the standard deviation of cosine similarity induced by this stochasticity. With the expected cosine similarity and its standard deviation per prompt, we computed a normalized Z-score for the observed image relative to its replication distribution. This Z-score measures the extent to which the observed image is better or worse than what’s expected for that prompt and is used for further analysis in section G.
@@
We generated these additional samples both for the original prompts on their assigned treatment arms and by replaying prompts on the counterfactual arms, as introduced in Figure 2 in the main text. Importantly, OpenAI updated its content filters between our initial experiment and image re-sampling. As a result, some prompts that originally produced images either generated no images or fewer images than requested during our regeneration attempts. This affected 1.8% (371 out of 18,990 prompts) of the data in our sample under the ”replaying” procedure (Section C.C.3.3).
Appendix E Main Text Analyses
This section provides detailed methodological information about the analyses presented in the main text.
E.1 Conceptual Framework
This appendix provides detailed derivations and proofs for the conceptual framework presented in the main text. While the main text introduces the conceptual framework and key insights, here we describe the model in more detail and develop its mathematical foundations and results more thoroughly.
We model a task where users attempt to replicate an object (such as an image) using a generative model. Perfect replication yields a maximum quality value of 1. The bounded quality function is defined as:
[…]
where $\theta\in(0,1]$ denotes the model capacity, $s\in(0,1]$ denotes the user’s skill at writing prompts, and $x\geq 0$ is the level of user’s effort in generating the appropriate prompts.
This function satisfies two intuitive properties:
  1. 1.
    Boundedness: Even as $x$ becomes large, $Q$ remains below 1, reflecting that the user cannot exceed perfect replication.
  2. 2.
    Diminishing returns: As user effort ($x$), increases, its marginal impact on quality becomes smaller.
We assume a linear cost function of prompting:
$c(x)\;=\;k\,x,$
where $k>0$ represents the marginal cost of user effort, for example in terms of time spent. The user’s utility given the quality and cost functions is
$U(\theta,s,x)=1-e^{-\theta sx}-k\,x.$
E.1.1 User Optimization
The user chooses the effort $x\geq 0$ such that it maximizes $U(\theta,s,x)$. We assume the user’s solution is always interior with positive effort which requires $\frac{\theta s}{k}>1$. This can be justified since the time cost associated with the task is minimal ($k\approx 0)$. As such, optimal user effort is $x^{*}(\theta,s)$:
$x^{*}(\theta,s)\;=\;\frac{1}{\theta s}\,\ln\Bigl{(}\frac{\theta s}{k}\Bigr{)}>0$
Thus, the optimal quality the user achieves at their optimal effort level becomes:
$Q^{*}=Q\big{(}\theta,s,x^{*}(\theta,s)\big{)}=1-\frac{k}{\theta s}>0$
E.1.2 Comparative Statics
We now analyze how quality, $Q^{*}$, varies with $\theta$, $s$, and $k$. Comparative statics are easily shown by taking the derivative of $Q^{*}$ with respect to the model capacity, user’s skill, and prompting cost.
Model Capacity: As model capacity $\theta$ increases, optimal quality $Q^{*}$ increases.
$\frac{\partial Q^{*}}{\partial\Theta}\Bigg{|}_{\Theta=\theta}=\frac{k}{\theta^{2}s}>0$ (11)
User’s Skill: As the skill of the user $s$ increases, the optimal quality $Q^{*}$ also increases.
$\frac{\partial Q^{*}}{\partial S}\Bigg{|}_{S=s}=\frac{k}{\theta s^{2}}>0$ (12)
Prompting Cost: As the prompting cost $k$ increases, the optimal quality $Q^{*}$ decreases.
$\frac{\partial Q^{*}}{\partial K}\Bigg{|}_{K=k}=-\frac{1}{\theta s}<0$ (13)
In the three derivatives above, the capital letters $\Theta,S$ and $K$ correspond to variables with respect to which derivatives are taken, and small letters $\theta,s$ and $k$ are the values at which derivatives are evaluated.
E.1.3 Model Improvement and Skill Heterogeneity
We next show that improvements in model capacity reduce the performance gap between high and low prompting skills. This result is mainly due to the diminishing returns in output quality.
Result. As model capacity $\theta$ improves, the effect of user skill $s$ on optimal quality $Q^{*}$ diminishes.
Proof: This can be easily shown by taking the derivative of equation 12, which denotes the total effect of user skill, with respect to a model improvement ($\theta$).
$\frac{\partial^{2}Q^{*}}{\partial S\partial\Theta}\Bigg{|}_{S=s,\Theta=\theta}=-\frac{k}{\theta^{2}s^{2}}<0$ (14)
As before capital letters $\Theta,S$ correspond to variables that are evaluated at $\theta,s$. $\square$
E.1.4 Decomposition into Model and Prompt Effects
As discussed in the main text, we decompose the total quality improvement into two parts: a model effect (the improvement from upgrading the model but holding the prompting effort fixed) and a prompting effect (the improvement from re-optimizing the effort in response to the upgraded model). To perform this decomposition, we consider a “counterfactual” quality function that uses one model’s capacity but another model’s optimal prompt:
$Q\bigl{(}\theta_{2},s,x^{*}(\theta_{1},s)\bigr{)}=1-\exp\big{(}-\theta_{2}s\,x^{*}(\theta_{1},s)\big{)}$ (15)
which is the quality obtained by using the model $\theta_{2}$ while choosing the prompting effort $x^{*}(\theta_{1},s)$ that would have been optimal if the user actually had model $\theta_{1}$. Note that if $\theta_{1}=\theta_{2}$, then $x^{*}(\theta_{1},s)$ is indeed the true optimum for $\theta_{2}$, yielding $Q^{*}$.
Given the formulation above, we can now express the model effect ($M$) and the prompting effect ($P$) by taking the derivative of the counterfactual quality in equation 15 with respect to the model capacity directly or indirectly through its effect on the prompt. The model effect is the derivative of 15 with respect to the direct effect of model capacity, while keeping the user’s effort fixed.
$M$ $=\;\frac{\partial Q\bigl{(}\Theta_{2},s,x^{*}(\Theta_{1},s)\bigr{)}}{\partial\Theta_{2}}\Bigg{|}_{\Theta_{1}=\Theta_{2}=\theta}$ (16)
$=\;s\cdot x^{*}(\Theta_{1},s)\cdot\exp\left(-\Theta_{2}s\cdot x^{*}(\Theta_{1},s)\right)\Bigg{|}_{\Theta_{1}=\Theta_{2}=\theta}$
$=\;\frac{k}{\theta^{2}s}\cdot\ln\left(\frac{\theta s}{k}\right)$
We can obtain the prompt effect by taking the derivative of 15 with respect to its indirect impact on the user’s effort, while keeping the direct impact of the model capacity fixed.
$P$ $=\;\frac{\partial Q\bigl{(}\Theta_{2},s,x^{*}(\Theta_{1},s)\bigr{)}}{\partial\Theta_{1}}\Bigg{|}_{\Theta_{1}=\Theta_{2}=\theta}$ (17)
$=\;\Theta_{2}s\cdot\exp\left(-\Theta_{2}s\cdot x^{*}(\Theta_{1},s)\right)\cdot\,\frac{\partial x^{*}(\Theta_{1},s)}{\partial\Theta_{1}}\Bigg{|}_{\Theta_{1}=\Theta_{2}=\theta}$
$=\;\frac{k}{\theta^{2}s}\left(1-\ln\left(\frac{\theta s}{k}\right)\right)$
We now show that the model effect mainly benefits low-skilled users in contrast to the prompting effect which mainly benefits the high-skilled users.
Proposition 3.
As user skill $s$ increases, the model effect becomes smaller, for sufficiently small prompting cost $k$.
Proof.
To examine how the model effect varies by user skill, we can take the derivative of the model effect M in equation 16 with respect to $s$.
$\frac{\partial M}{\partial S}\Bigg{|}_{S=s}$ $=\frac{\partial}{\partial S}\frac{k}{\theta^{2}S}\cdot\ln\left(\frac{\theta S}{k}\right)\Bigg{|}_{S=s}$
$=\frac{k}{\theta^{2}s^{2}}\left(1-\ln\left(\frac{\theta s}{k}\right)\right)<0$
for sufficiently small $k$. ∎
Intuitively when prompting costs are low, users can afford to put substantial effort into writing prompts (so $x^{*}$ is high). Once the effort is large enough, the marginal effect of the model on quality (i.e. the model effect) will be decreasing by user skill. This happens due to the diminishing returns in the quality function. Once a high-skill user is already achieving near-peak quality by putting in large effort, small increases in $\theta$ deliver small additional benefit. However, the same model improvement yields a much larger marginal impact for a low-skilled user who is putting the same level of effort.
Proposition 4.
As user skill $s$ improves, the prompting effect becomes larger, for sufficiently small cost $k$.
Proof.
Proof To examine how the prompt effect varies by user skill, we can again take the derivative of the prompting effect P in equation 17 with respect to $s$.
$\frac{\partial P}{\partial S}\Bigg{|}_{S=s}$ $=\frac{\partial}{\partial S}\frac{k}{\theta^{2}S}\cdot\big{(}1-\ln(\frac{\theta S}{k})\big{)}\Bigg{|}_{S=s}$ (18)
$=\frac{k}{\theta^{2}s^{2}}\left(\ln\left(\frac{\theta s}{k}\right)-2\right)>0$
for sufficiently small $k$. ∎
Comparing equations E.1.4 and 18, we conclude that the prompting effect is less sensitive to user skill than the model effect, as its marginal with respect to skill has a smaller magnitude. This makes the overall marginal effect of user skill negative, matching the earlier finding in equation 14.
E.2 Task Performance and ATEs